I want to produce 200 pounds of mixed nuts which I can sell for $4.45 per pound. The mix uses peanuts at $2.19 per pound and almonds at $6.47 per pound. How many pounds of both peanuts and almonds are needed to create the perfect mix?
Found 2 solutions by AnimalLover, stanbon:
Answer by AnimalLover(6)

So, you want less almonds since they cost more...
And that would mean more peanuts to keep it equal...
125 pounds of peanuts and 75 pounds of almonds.

Answer by stanbon(75887)

I want to produce 200 pounds of mixed nuts which I can sell for $4.45 per pound. The mix uses peanuts at $2.19 per pound and almonds at $6.47 per pound. How many pounds of both peanuts and almonds are needed to create the perfect mix?
Equation:
value + value = value
----
2.19x + 6.47(200-x) = 4.45*200
--------------
219x + 647*200 - 647x = 445*200
------
-428x = -202*200
-----------
x = 94.39 lbs (amt of peanuts needed)
----------
200-94.39 = 105.61 lbs (amt of almonds needed)
-------
Cheers,
Stan H.
-----------